OpenStudy (anonymous):

simplify: 1/cot(x)-1/2sinxsecx

OpenStudy (anonymous):

algebra is easier i you replace \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\)

OpenStudy (anonymous):

\[\frac{1}{\frac{a}{b}}-\frac{1}{2b\times \frac{1}{a}}\]

OpenStudy (anonymous):

you get \[\frac{b}{a}-\frac{a}{2b}\]

OpenStudy (anonymous):

this isn't algebra it's trigonometry. I don't understand what you are doing.

OpenStudy (anonymous):

then subtract, get \[\frac{2b^2-a^2}{2ab}\]

OpenStudy (anonymous):

it is algebra, just with trig functions

OpenStudy (anonymous):

is the question \[\frac{1}{\cot(x)}-\frac{1}{2\sin(x)\sec(x)}\]?

OpenStudy (anonymous):

yes, but i need to answer it using trig identies.

OpenStudy (anonymous):

the reciprocal of cotangent is tangent, and the reciprocal of secant is cosine

OpenStudy (anonymous):

so you can write \[\tan(x)-\frac{\cos(x)}{2\sin(x)}\] if you like

OpenStudy (anonymous):

is the answer tanx

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you can subtract if you like

OpenStudy (anonymous):

but the answer is not some simple trig function, at least not to the question you posted there is a 2 in the denominator right?